OpenStudy (anonymous):

Which is an equation of the line passing through points at (1, 6) and (-2, 1)? A. y = 2x + 5 B. y = 5/3x + 13/3 C. y = x + 5 D. y = -1/2x +13/2

Parth (parthkohli):

Find the slope first. \(\Large \color{MidnightBlue}{\Rightarrow {1-6 \over -2 - 1} }\) \(\Large \color{MidnightBlue}{\Rightarrow {-5 \over -3} }\) \(\Large \color{MidnightBlue}{\Rightarrow {5 \over 3} }\) Till now we know that the equation is y = (5/3)x + ? Let's find the y intercept. y intercept would be 13/3.

OpenStudy (anonymous):

the standard equation is y=mx+c. m=slope given as (y2-y1)/(x2-x1). So in this problem m= (1-6)/(-2-1)=-5/-3=5/3. and from either given points y = mx+c you will get c = 13/3. so B is the answer.

OpenStudy (anonymous):

Thanks !